Science Report – March 6th

Crew 263 Scientific Report 06-03-2022

SOL: 13

Name of person filing report: Léa Rouverand

Geology Field Campaign:

Geology field campaign by human operators has been investigated throughout the mission second week (SOL 7 to SOL 13). Analysis using a SciAps Z-903 handheld LIBS (laser-induced breakdown spectroscopy) analyzer has been performed as well as outcrop sketching, image capture and sample collection. The participants are now used to performing such tasks under EVA protocol and spacesuits which constraints their movement and field vision. The study has put into light the large flexibility of human operators to perform field study: 5 outcrops were analyzed in a total of 3 outings, the EVA participants could move across units and perform studies at different elevations. On-field decision making was also recorded by using a drone to chose on the field the outcrops to analyze. The J2/K1 transition was identified in terms of rock textures and gypsum were found chemically using the LIBS.


This analysis uses Aquapads to test the drinking water quality. The water was found to be very clean and safe for consumption.

Ultrasound surveillance:

On SOL 10, image capture of 5 organs for each Crew Members were performed using an ultrasound device and an Augmented Reality software. It is the same experiment performed on SOL 2. However, an improvement on the ultrasound image capture performance by the Crew Members have been observed.

HF wave propagation experiment:

This experiment aims to study the propagation characteristics of the ionosphere for HF waves to achieve contact between MDRS and Toulouse (France). On SOL 10, we successfully captured HF signals on 2 different frequencies (21 MHz and 28 MHz). Unfortunately, they did not hear back from us. Another trial will be performed 17.


This experiment aims to study the effect of gravity on plant growth. Lentil and bean seeds are planted on a rotating platform. The seeds suffered from the relocation from the GreenHab to the ScienceDome and/or local atmosphere. We have placed them in water to wait for initial growth before planting them back in the rotating platform.

Drone Astronaut Rescue:

The experiment’s aim is to assess the use of an ANAFI thermal drone for Astronaut Rescue. A pair of 2 astronauts goes away from the other pair. The efficiency of a time lee identification of the “hidden” pair using a drone is to be investigated. First trial was performed on SOL 4 and a second trial was performed on SOL 10, the drone allowed to spot the Astronaut pair in 5 minutes (longer than on SOL4 as the “hidden” participants were in the shadow) and join them in 7 minutes (shorter than on SOL 4, as the drone was left at the target location and not turned back to the drone operator, real-time feedback was possible). The use of the drone thermal feature is still to be investigated as it still made the image more confusing. We also need to compare the rescue protocol to one without a drone; this is planned on SOL 16.

On SOL 10, the “hidden” pair also investigated emergency protocols. The protocol is made by middle school students and serve as an Educational Activity to introduce them to crewed space missions and the scientific methodology. One participant simulated an arm injury, the other participant had to follow the emergency protocol to handle the injured participant. A stress analysis was also performed at the same time as the simulation injured participant was not the initially planned person. One participant during the EVA and in the Station was aware of this “switch” in person. Feedbacks has been obtained on the protocol.

Hab depressurization emergency protocol:

On SOL 11, a HAB depressurization emergency protocol, also made by middle-school students, was tested out. An initial disorganization of the crew members in following the protocol was observed but they successfully managed to follow the protocol in time and did not lead to “simulated” fatalities.

Atmospheric Experiment:

There is a total of 5 atmospheric experiments; LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter) and a weather station. Mega-Ares has been out because of an assembly issue but has been fixed and reinstallation was performed on SOL 6. The Electrical Field Mill does not turn on so is back at the station until we have researchers’ feedback.

Augmented Reality Testing:

First trials of the Augmented Reality Application testing were performed between SOL 1 and SOL 4. Software issues with menu placement and hardware issues were identified.

On-going experiments: Sleep study using Dreem Headbands, TELEOP (analysis of spatial awareness and isolation on teleoperation tasks), Psychosocial and Cognitive questionnaires, Psychometric tests, nothing to declare on SOL 13.

Questions, concerns and requests to Mission Support: NaN

Journalist Report – March 6th

Sol 13 – Sweet dreams

Author: Nicolas Wattelle

Against all odds, this morning’s weather has been wonderful. A nice and warm sun was waiting for us to wake up. The wind was slightly blowing, but nothing compared to what was announced. Mars is full of surprises.

So, we were allowed to go out and do the geology outing we could not do the day before, we even took advantage of it to realise maintenance on the HF antenna, change the outdoor batteries, and put the LOAC back on the field (which we took back in to protect it from the potential snowfall). So, at the end of the day, it was a big EVA, especially since the geology study needed us to go to a place named “White moon”, the furthest away we have gone so far. Hence, we had a great breakfast, carefully packed our needed material, and left the Hab for this morning adventure.

We were four, Cerise, Marine, Léa and I, heading to the North, and it went well. We were happy for the geology analysis we made there, and for the new landscapes we discovered…

It did not look like a “classic” Sunday of a Supaero Crew, it was more like a mix between a rest day and a work day (because we already took some rest the day before). Elena went through data analysis of the Teleop experiment (the one where martionauts teleoperate rovers in different positions) and cooked, Valentine tided up the lower deck and spent time drawing, Cerise went on EVA this morning, adjusted mission planning and knitted during the afternoon… To sum up, we worked in a relaxed atmosphere!

Most of us had a (needed) restorative night, the awakening was sweeter than the other days. And our numbers speak for themselves: we have a complete follow-up of our nights using Dreem headbands. These devices are really easy to use and are efficient in the context of our project! We just need to put them before going to sleep (like Cerise does on the photo), and they analyse our brain waves, breath rhythm, heartbeat, and even our detailed movements through the night (you would be impressed discovering how much a person moves during a night…)!

They allow us an access to durations and proportions of sleeping phases: deep, light, rapid eye movement (REM)… What we are interested in, is how all of this data evolves throughout a mission like ours, and also to keep an eye on it and take measures (reorganisation of the planning, relaxation exercises, naps…) if needed. Their data is thus very useful to efficiently keep our mission running.

You might ask yourself if it is possible to sleep correctly with it: do not worry, they are not very invasive, after two or three nights of accommodation (that we did weeks before the mission) you almost don’t feel it anymore. You can peacefully have sweet dreams!

EVA Report – March 6th

Crew 263 EVA Report 06-03-2022

EVA #17

Author: Cerise Cuny

Purpose of EVA: Re-orienting the HF antenna, Changing the batteries of atmospheric experiments, Looking for gypsum for analysis along cow dung road, AR application testing for geological studies.

Start time: 9:04

End time: 11:30

Narrative: Before going out, we collectively decided to swap the work around the antenna and the battery change ; it seemed so much more logical to start with the antenna instead of going to the atmospheric experiment site, than back to the Hab and then go to the north. After checking the rovers usage and batteries, we thus headed to the antenna, it was a quick and easy task. We left the Hab with the rovers at 9:22 with Curiosity and Opportunity (both were at 100% battery and hour usages were 169.2 and 85.2, respectively). After a quick battery change, we left the atmospheric experiment site at 9:32. We successfully spotted the pole indicating the intersection was a hundred meters further. We then struggled a bit to find the road. When it was obvious we went too far (200m from the pole), we turned back and then found the road. We parked near the intersection at 10:00, (Batteries: 68% for Curiosity, 75% for Opportunity). We walked for an hour, stopping at several interesting spots that might have been relevant for the geological study. Léa and Marine (responsible for the geological experiment), did quick drawings, wrote down the GPS information and did LIBS analysis while Nicolas and I took pictures of the areas of interest to add to their analysis report. Analysis on the field did not reveal any gypsum but the samples we brought back to the Hab were indeed gypsum! The landscape was absolutely breathtaking, once more very different from what we were used too until now. We went back to the rovers at 10:56. On the way back, before Tank Wash, we spotted an outcrop that might have been relevant, stopped at 10:59 to check it out, and went back to the rovers again at 11:10. We reached the Hab, plugged the rovers in and then entered the Air Lock at 11:24. Curiosity had 61% battery charge and 169.9 usage, and Opportunity 65% and 85.8.

Destination: Around Hab (for antenna), Intersection between Entrance Road and Cow Dung Road (for atmospheric experiment), Intersection between Cow Dung Road and Brahe Hwy (for Gypsum search).

Coordinates (use UTM NAD27 CONUS): Atmospheric Experiment: 518288 E, 4250586 N ; Gypsum Search: 517700 E ,4254700 N

Participants: Cerise Cuny, Nicolas Wattelle, Léa Rouverand, Marine Prunier.

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road.

Mode of travel: Foot to go to the antenna, rovers to go to the atmospheric experiments, then to the area of interest and go back to the Hab.

Operations Report – March 6th

Crew 263 Operations Report 06-03-2022

SOL: 13

Name of person filing report: Mathéo Fouchet

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report


Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Opportunity rover used: Yes

Hours: 85.8

Beginning charge: 100%

Ending charge: 65%

Currently charging: Yes

Curiosity rover used: No

Hours: 169.9

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Perseverance rover used: No

Hours: 228.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

General notes and comments: Nothing to report

Summary of Hab operations:

WATER USE: 24 gallons

Water (static tank): 301 gallons

Water (loft tank): 31 gallons

Water Meter: 01562997 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: Yes

Notes: At 6pm we had a smell of gas coming from the vents in the bedroom and the one in the lower deck, with the advice of Shannon we used the two gas detector from the RAM and everything seems to be back to normal. We are going to calibrate and check agin with the gas detector every hour just in case.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations:

Dual split: Off

SOC: 89% at 5:45pm

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: See HSO Report

Questions, concerns and requests to Mission Support: Nothing to report

Sol Summary – March 6th

Crew 263 Sol Summary 06-03-2022

Sol: 13

Summary Title: Gypsum and Suspicious Gas Smell

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: Yesterday, we decided to start the sport session 30 minutes later than usual to have 30 minutes of bonus sleep. We will probably keep this timing for the rest of the mission as it made us feel more energetic. We were happy to see that there was no snow this morning, allowing us to do our EVA as planned, in the North, between Tank Wash and White Moons. We were not expecting gypsum as we saw it, but took samples to analyse in the Hab. After talking with Shannon about it and a quick LIBS analysis, we found out they were indeed Gypsum, which was quite satisfying. It was our first time in this area and we found it beautiful.

In the afternoon, there wasn’t a lot of scientific activities. We thus had to cheat boredom: some of us did colour painting supervised by our specialist, Marine, Nicolas and Mathéo did a video tour of the station, and I knitted. We had a little scare in the evening: a very strong smell of gas spread in the Hab, coming from the vents. Shannon radio-assisted us in the search for a leak: none was detected and 15 minutes later we could not smell it anymore.

Look Ahead Plan: We were very excited to plan Tuesday’s EVA that will probably be the most relevant one for the research team back in Toulouse that trained Marine and Lea to Geology and LIBS analysis. As it will be a very long and surely exhausting EVA, we planned tomorrow’s EVA accordingly, so that Tuesday’s EVA team rests the day before. Tomorrow I’ll turn 25: I am sure the Crew will make it a special day!

Anomalies in work: Gas odour in the Hab in the evening. No leak detected, and odour stopped.

Weather: Sunny in the morning, cloudy and very windy in the afternoon.

Crew Physical Status: Good.

EVA: EVA 17 in the morning. No issue to report.

Reports to be filed: Sol Summary, EVA Report (n°17), EVA Request (n°18), Operations Report, Journalist Report, HSO Report, Scientist Report.

Support Requested: No support requested.

Copyright © The Mars Society. All rights reserved. | Main Site